For a recent job, a plumber earned $28 per hour, adn the plumber's apprentice earned $15 per hour. The plumber worked 3 hours more htan the apprentice. If together they were paid $213, how much did each earn?
All I know is that I need to find an expression first for the number of hours each worked on the job, which would be something like (28x+3)+ 15x=213 ???
I'm really not sure, please help!
Thank you!

Found 2 solutions by ankor@dixie-net.com, Earlsdon:
Answer by ankor@dixie-net.com(12699)

For a recent job, a plumber earned $28 per hour, adn the plumber's apprentice earned $15 per hour. The plumber worked 3 hours more htan the apprentice. If together they were paid $213, how much did each earn?
:
All I know is that I need to find an expression first for the number of hours each worked on the job, which would be something like (28x+3)+ 15x = 213 ???
:
you're almost right (bracket), just solve for x to find the apprentice hrs
28(x+3) + 15x = 213
28x + 84 + 15x = 213
28x + 15x = 213 - 84
43x = 129
x = 129/43
x = 3 hrs
:
Find the wages:
Plumber: 6 * 28 = 168
Apprent: 3 * 15 = 45
-----------------------
our given total = 213,

Answer by Earlsdon(6103)

You've got the right idea, but...
Let x = the number of hours worked by the apprentice.
The plumber worked 3 hours more than that, so the plumber workrked (x+3) hours.
The plumber gets paid $28 per hour so he/she earns (x+3)($28) while the apprentice gets $15 per hour so he/she earns x($15). The sum of these earnings is $213. So here's the equation you need to solve for x:
(x+3)($28) + x($15) = $213 Simplify and solve for x.
28x + 84 + 15x = 213 Combine like-terms.
43x + 84 = 213 Subtract 84 from both sides.
43x = 129 Divide by 43
x = 3 This is the number of hours worked by the apprentice, and, at $15 per hour, he/she earns $45.
The plumber works x+3 hours or 6 hours, and, at $28 per hour, earns $168